### v2

@@ 4 Kernel models are conjunction-wise additive
Our primary theoretical contribution in this work is to characterize the full range of compositional computations that can be implemented by kernel models with disentangled inputs. We find that even though these models can learn arbitrary training sets, their test set behavior is restricted to adding up values assigned to each conjunction (combination of components) seen during training. We call this motif “conjunction-wise additivity.” Below, we formally state our finding and explain its implications.
@@ 4.1 Random networks yield compositionally structured representations
We first note that a linear readout of the disentangled input representation effectively constrains the model to adding up a value for each component (“component-wise additivity”):
$f(z)=w^{T}z=w^{T}\sum_{c=1}^{C}z_{c}=\sum_{c=1}^{C}f_{c}(z_{c}),\quad f_{c}(z_{c}):=w^{T}z_{c}.$ (2)
Although models of this class perfectly generalize on component-wise additive tasks, such as symbolic addition, they are incapable of even learning the training data (much less generalizing) on tasks that are not component-wise additive. In particular, these models cannot learn context-dependent computations or equivalence relations — both fundamental instances of compositional reasoning.
The nonlinear transform $\phi(z)$ can overcome this constraint; in particular, multi-layer nonlinear neural networks can learn arbitrary training data (in the infinite-width limit) (Hornik et al., 1989; Cybenko, 1989; Rigotti et al., 2013). In general, the wide range of possible neural network architectures and their nonlinearities makes it difficult to derive general statements about their representations, but in this case we can take advantage of a particularly useful property of disentangled categorical inputs. That is, its induced kernel, $K(z,z^{\prime})=z^{T}z^{\prime}$, depends only on the number of components that are identical between both inputs (specifically, $K(z,z^{\prime})=c$ if $c$ components are identical). We call any representation that satisfies this criterion “compositionally structured” (see Section A.2 for further motivation). Notably, in the infinite-width limit, the kernel of random weight neural networks, $K_{\phi}(z,z^{\prime})=\phi(z)^{T}\phi(z^{\prime})$, depends only on the input kernel $K(z,z^{\prime})$, making $K_{\phi}(z,z^{\prime})$ also compositionally structured (Cho & Saul, 2009, see Section A.3). This representational constraint allows us to derive strong constraints on the models’ generalization behavior.
@@ 4.2 Compositionally structured kernel models are conjunction-wise additive
$\text{Conj}(z|Z^{\text{train}}):=\left\{J\subseteq\{1,\dotsc,C\}|\exists z^{\text{tr}}\in Z^{\text{train}}:\forall c\in J:z_{c}=z^{\text{tr}}_{c}\right\}.$ (3)
We then prove:
@@
Consider a kernel model $f$ with a compositionally structured representation and training data $Z^{\text{train}}$. For any such model, we can find conjunction-wise functions $f_{J}:\prod_{c\in J}Z_{c}\to\mathbb{R}$, where $J\subseteq\{1,\dotsc,C\}$, such that for any input $z\in Z$, the model response is given by
$f(z)=\sum_{J\in\text{Conj}(z|Z^{\text{train}})}f_{J}(z_{J}),\quad z_{J}:=(z_{c})_{c\in J}.$ (4)
We prove Theorem 4.1 in Section A.4, but to develop an intuition for why this theorem holds, we consider an example representation concatenating one-hot vectors representing each component and conjunction of components. For example, if $C=2$, this representation will concatenate one-hot representations of each component with a one-hot representation of the conjunction of both components. The readout weight for a conjunction $z_{J}$ seen during training will change (in Eq. 4, this weight is given by $f_{J}(z_{J})$). In contrast, the weights associated with conjunctions not seen during training remain at their initial value of zero (which is why we only sum over $J\in\text{Conj}(z|Z^{\text{train}})$).
@@ 4.3 Conjunction-wise additivity constrains the tasks kernel models can solve
Theorem 4.1 immediately implies that compositionally structured kernel models cannot solve any task that cannot be expressed in conjunction-wise additive terms. This highlights fundamental computational restrictions on a broad range of models — including pre-trained models with disentangled representations that are finetuned in the kernel regime and infinite-width neural networks. Intriguingly, these restrictions are not caused by an architectural constraint, as the model can learn arbitrary training data. Rather, we highlight restrictions on how models – without any architectural constraints – can generalize (Zhang et al., 2021). Below we spell out the consequences of these restrictions.
@@
First, for inputs with two components, the model’s behavior on the training set can be expressed as $f(z)=f_{1}(z_{1})+f_{2}(z_{2})+f_{12}(z)$. However, on the testing set, $f_{12}(z)$ falls away and the model is constrained to be component-wise additive: $f(z)=f_{1}(z_{1})+f_{2}(z_{2})$. A striking consequence of this is that compositionally structured models – including infinite-width neural networks and regardless of implementational details – can only generalize on component-wise additive tasks. In particular, this implies that these models cannot generalize on transitive equivalence — a fundamental instance of relational reasoning. Intriguingly, transitive ordering – a superficially similar task – can be solved by a kernel model. This highlights the importance of a formal perspective on compositional tasks.
@@
For inputs with more than two components, a conjunction-wise additive computation additionally encodes partial conjunctions between components (e.g. 12, 13, and 23 for three components). For test inputs, the model cannot rely on the unseen full conjunction 123, but it can rely on any partial conjunction it has seen before. To see whether a kernel model can, in principle, generalize on such a task, we must 1) identify the conjunctions seen during training, $\text{Conj}(z|Z^{\text{train}})$, for each $z\in Z^{\text{test}}$, and 2) determine whether the target can be written as a conjunction-wise sum. For example, for context dependence, for all new test inputs, we have seen the context-feature conjunctions 12 and 13, but not the feature-feature conjunction 23. As a result, model behavior on the test set can be written as
$f(z)=f_{1}(z_{\text{co}})+f_{2}(z_{f1})+f_{3}(z_{f2})+f_{12}(z_{\text{co}},z_{f1})+f_{13}(z_{\text{co}},z_{f2}).$
While context dependence is not component-wise additive, it is conjunction-wise additive. This is because we can encode it using $f_{12}$ and $f_{13}$: $f_{12}(z_{\text{co}},z_{f1})$ encodes the target when the context $z_{\text{co}}$ indicates the $z_{f1}$ as relevant and is zero otherwise; $f_{13}(z_{\text{co}},z_{f2})$ works in the opposite way. Thus, the model can, in principle, identify a set of weights that would allow it to generalize correctly.
This example illustrates that conjunction-wise additivity tells us both whether kernel models with compositionally structured representations can solve a certain task, and also how they solve it. As we noted, these insights immediately transfer to any model with compositionally structured representations that is trained or fine-tuned in the kernel regime. Because our theory determines the compositional computations emerging from dataset statistics, it is also likely to shed light on other learning models, including feature-learning deep neural networks (see Section 6).
@@
Lastly, our theory tells us how to make a task non-additive: in Section D.2.3, we describe such a modification to context dependence that tests generalization to novel context-feature conjunctions. By helping us design hard benchmark tasks that are not solvable by kernel models, our framework grounds research into learning mechanisms that can implement other compositional computations.
@@
Conjunction-wise additivity only determines whether kernel models can, in principle, solve a certain task. However, model generalization also depends on whether the model identifies the correct conjunction-wise function. This depends on the model’s representational geometry and the dataset statistics. As noted in Section 2, the behavior of kernel models depends on their representation $\phi(z)$ only through the induced kernel $K_{\phi}(z,z^{\prime})$. In this section, we employ the modular approach suggested by this insight: We first investigate how different choices in network architecture influence its induced kernel (Section 5.1). Then, we characterize task performance on symbolic addition and context dependence across the full range of compositionally structured representations (Section 5.2).
@@ 5.1 Overlap salience characterizes compositional representational geometry
In Section 4, we found that a broad range of transformations yield compositionally structured representations, i.e. representations that have the same similarity $K(z,z^{\prime})$ for all inputs $z,z^{\prime}$ overlapping in the same number of components. To characterize representations in this space, we introduce a new metric: representational salience. We formally define this metric in Section B.1. Intuitively, for $c=1,\dotsc,C$, the salience $S(c;C)$ measures the unique contribution of the subpopulation representing a conjunction of $c$ components. Further, $S(c;C)$ is normalized so that the salience of all components and conjunctions of components sums up to one. For example, the input representation only encodes single components and therefore $S(1;C)=1/C$ and, for all $c>1$, $S(c;C)=0$.
$S(c;C)$ is computed from the representational similarities $K(z,z^{\prime})$. It is useful because unlike those similarities, it makes immediately apparent how strongly different conjunctions are encoded (Section B.2). Further, it reduces the set of $C+1$ similarities characterizing the space of compositionally structured representations to $C-1$ free parameters: It removes the overall magnitude of the representation (by normalizing) and the baseline activity (by excluding the similarity between inputs with no overlap). This choice reflects the fact that both of these parameters often have negligible impact on model behavior (Section B.3).
We now analyze how $S(c;C)$ evolves over different layers of a random network, using an input with three components as an example (Fig. 2; see Section B.4 for other values of $C$). At the input stage, only individual components are represented and so $S(2;3)$ and $S(3;3)$ are zero. As the network gets deeper, these saliences increase, whereas $S(1;3)$ decreases. Eventually, the salience of the intermediate conjunction $S(2;3)$ decreases again, whereas $S(3;3)$ continues to increase. Indeed, for ReLU networks, we prove the following statement:
Proposition 5.1.
We prove the proposition in Section B.5. It implies that in very deep networks, the full conjunction dominates the representation, making compositional generalization to overlapping inputs harder.
@@ 5.2 Kernel models suffer from memorization leak and shortcut bias
While there are many readout weights that can lead to minimal error, kernel models trained with gradient descent or ridge regression learn the weights with minimal $\ell_{2}$-norm (see Section 2). We now characterize how this inductive bias influences compositional generalization across different representational geometries. To characterize the full range of compositionally structured representations, we compute the behavior of the model directly using the kernel (see Appendix C). Importantly, as implied by our modular approach, our analysis clarifies the behavior of all random weight neural networks, including those covered in the previous section.
Symbolic addition suffers from a memorization leak. We first consider an example case ($S(1;2)=0.4,\mathcal{W}=\{0\}$). We find that though the model perfectly learns the training cases, it underestimates the test cases by a proportional factor (Fig. 3a). To understand why, we consider the model’s functional form on the training cases: $f(z)=f_{1}(z_{1})+f_{2}(z_{2})+f_{12}(z)$ (see Section 4). Intuitively, $\ell_{2}$-norm minimization tends to yield distributed weights. As a result, unless the salience of the conjunction $S(2;2)$ is zero, the kernel model will associate some non-zero weight with it (i.e. $f_{12}(z)\neq 0$). This, in turn, necessarily distorts the generalization on the test set, $f(z)=f_{1}(z_{1})+f_{2}(z_{2})$.
We call this tendency to use the full conjunction (and its effect on generalization) “memorization leak.” To characterize it more systematically (and add mathematical rigor to the intuition above), we analytically characterize model behavior on a general family of symbolic addition tasks:
@@
We prove the proposition in Section D.1. It implies that for this entire task family, model generalization is distorted by a proportional factor $m$ (see Fig. 3b). The formula implies that $m<1$ as long as $S(1;2)<\tfrac{1}{2}$ (i.e. as long as $S(2;2)>0$). Further, $m$ is smaller for lower $S(1;2)$ and smaller training set size $p$. (Note that $\mathcal{W}$ corresponds to the rows/columns making up the example training sets in Fig. 1b.) Interestingly, these are the only two factors that influence $m$. In particular, while interpolation is often seen as easier than extrapolation, this does not impact model behavior here.
More broadly, the memorization leak is a ubiquitous issue for statistical learning models trained on compositional tasks: $\ell_{2}$-norm minimization tends to yield distributed weights and so if the conjunctive population is represented, the model will generally end up partially relying on this conjunction. As shown above, this necessarily distorts generalization (see also Section D.3 for a minimal example).
@@
Context dependence suffers from a shortcut bias. Next, we empirically analyzed model generalization on context dependence across different representational geometries and training sets. We found that on a given task, each model either generalizes with 100% or 0% accuracy. For CD-3 (the task version leaving out the largest subset of feature combinations), the model only generalizes when $S(2;3)$ is high relative to $S(1;3)$ (Fig. 3c, dark green area). As a result, whether the network generalizes is highly sensitive to the nonlinearity and depth of the network. In contrast, for CD-2 and CD-1, a much wider range of representational geometries generalizes successfully.
To understand this phenomenon, we determined the total magnitude of model weights associated with the different conjunctions (Fig. 3d). We found that unsuccessful models (e.g. blue color in Fig. 3d) had much larger magnitudes associated with the context component and the full conjunction. Notably, on CD-3, context is highly correlated with the target and can predict 2/3 of the training data. Models with high $S(1;3)$ (context) and $S(3;3)$ (full conjunction) exploit the context shortcut and use the full conjunction to learn the remaining training data. This strategy explains why these models fail to generalize to the test set. For CD-2, context is much less predictive on the training data (accuracy of $\tfrac{9}{16}$). This explains why only very low $S(2;3)$ yields failure to generalize on CD-2 or CD-1.
Our analysis shows how model and task structure interact to either give rise to a generalizable rule or a statistical shortcut (see Section D.3 for a minimal example). This highlights that for compositional tasks with strong spurious correlations, model behavior will be highly sensitive to architectural details affecting the representational geometry such as depth and nonlinearity.
6 Our theory describes the behavior of deep networks on conjunction-wise additive tasks

### v3

@@ 4 Kernel models are conjunction-wise additive
Our primary theoretical contribution is to characterize the full range of compositional computations that can be implemented by kernel models with compositionally structured inputs. We find that even though these models can learn arbitrary training sets, their test set behavior is restricted to adding up values for each conjunction (combination of components) seen during training. We call this motif “conjunction-wise additivity.” Below, we formally state our finding and explain its implications.
@@ 4.1 Random networks yield compositionally structured representations
We first note that a linear readout of the multi-hot input constrains the model to adding up a value for each component (“component-wise additivity”): $f(x)=\sum_{c=1}^{C}f_{c}(z_{c})$. Although models of this class perfectly generalize on component-wise additive tasks, such as symbolic addition, they are incapable of even learning the training data (much less generalizing) on tasks that are not component-wise additive. In particular, these models cannot learn context-dependent computations or equivalence relations — both fundamental instances of compositional reasoning.
The nonlinear transform $\phi(x)$ can overcome this constraint; in particular, multi-layer nonlinear neural networks can learn arbitrary training data (in the infinite-width limit) (Hornik et al., 1989; Cybenko, 1989; Rigotti et al., 2013). Normally, the wide range of possible network architectures makes it difficult to derive general statements about their representations, but in this case we can take advantage of the fact that $\phi(x)$ will also be compositionally structured (Fig. 1c):
Proposition 4.1.
For a random weights neural network $\phi$ with a compositionally structured input $x$, in the infinite-width limit, $\phi(x)$ is also compositionally structured.
Proposition 4.1 holds because in the infinite-width limit, the kernel of random weight neural networks depends only on the input kernel (Cho & Saul, 2009, see Section A.3). We now consider the constraints that compositional structure imposes on the models’ generalization behavior.
@@ 4.2 Compositionally structured kernel models are conjunction-wise additive
$\text{Conj}(z|Z^{\text{train}}):=\left\{J\subseteq\{1,\dotsc,C\}|\mbox{there is }z^{\text{tr}}\in Z^{\text{train}}\mbox{ such that }z_{c}=z_{c}^{\text{tr}}\mbox{ for all }c\in J\right\}.$ (1)
@@
For any kernel model $f$ with a compositionally structured representation, we can find conjunction-wise functions $f_{J}:\prod_{c\in J}Z_{c}\to\mathbb{R}$, where $J\subseteq\{1,\dotsc,C\}$, such that for any input $x\in\mathbb{R}^{d}$ representing components $z\in Z$, the model response is given by
$f(x)=\sum_{J\in\text{Conj}(z|Z^{\text{train}})}f_{J}(z_{J}),\quad z_{J}:=(z_{c})_{c\in J}.$ (2)
We prove the theorem in Section A.4. It implies that for a given test input, kernel models add up a value for each component and partial conjunction seen during training (Fig. 1d). We call this computation “conjunction-wise additive.” In particular, for inputs with two components, the model’s behavior on the training set can be expressed as $f(x)=f_{1}(z_{1})+f_{2}(z_{2})+f_{12}(z_{1},z_{2})$. Thus, $f_{12}(z_{1},z_{2})$ enables the model to learn arbitrary training data. However, if $x$ represents $z\in Z^{\text{test}}$, the training set does not contain any input with the conjunction $(z_{1},z_{2})$ and the model is component-wise additive: $f(x)=f_{1}(z_{1})+f_{2}(z_{2})$. For inputs with more than two components, a conjunction-wise additive computation additionally encodes partial conjunctions seen during training, e.g., if $Z^{\text{train}}$ contains the conjunction $(z_{1},z_{2})$, $f(x)=f_{1}(z_{1})+f_{2}(z_{2})+f_{3}(z_{3})+f_{12}(z_{1},z_{2})$.
@@ 4.3 Conjunction-wise additivity constrains the tasks kernel models can solve
Theorem 4.2 implies that compositionally structured kernel models can only solve tasks that can be expressed in conjunction-wise additive terms. This highlights a fundamental computational restriction, partitioning the compositional task space into tasks that can and cannot be solved by kernel models (Fig. 2d). Intriguingly, these restrictions are not caused by an architectural constraint, as the model can learn arbitrary training data (at least with a nonlinear transformation $\phi$, see Section 5.1). Rather, we highlight restrictions on how models – without any architectural constraints – can generalize (Zhang et al., 2021). We now spell out the consequences of the highlighted restrictions.
@@
First, for inputs with two components, we noted above that the model’s behavior on the test set is component-wise additive. A consequence of this is that compositionally structured models can only generalize on component-wise additive tasks. In particular, this implies that these models cannot generalize on transitive equivalence — a fundamental instance of relational reasoning. Intriguingly, transitive ordering – a superficially similar task – can be solved by a kernel model. This highlights the importance of a formal perspective on compositional tasks.
@@
More generally, to see whether a kernel model can, in principle, generalize on a task with more than two input components, we must 1) identify, for each $z\in Z^{\text{test}}$, the conjunctions seen during training, $\text{Conj}(z|Z^{\text{train}})$, and 2) determine whether the target can be written as a conjunction-wise sum (Eq. 2). For context dependence for example, for all $z\in Z^{\text{test}}$, we have seen the context-feature conjunctions 12 and 13, but not the feature-feature conjunction 23. As a result, model behavior on a test trial $x$ representing the components $z_{\text{co}},z_{f1},z_{f2}$ can be written as $f(x)=f_{1}(z_{\text{co}})+f_{2}(z_{f1})+f_{3}(z_{f2})+f_{12}(z_{\text{co}},z_{f1})+f_{13}(z_{\text{co}},z_{f2}).$ Importantly, this conjunction-wise additive function can encode context dependence, specifically with the following schema: $f_{12}(z_{\text{co}},z_{f1})$ encodes the target when the context $z_{\text{co}}$ indicates $z_{f1}$ as relevant and is zero otherwise; $f_{13}(z_{\text{co}},z_{f2})$ works in the opposite way. Thus, the model can, in principle, identify a set of weights that would allow it to generalize correctly.
@@
This example illustrates that conjunction-wise additivity tells us both whether kernel models with compositionally structured representations can solve a certain task, and also how they solve it. It even tells us how to make a task non-additive: in Section E.2.3, we describe such a modification to context dependence that tests generalization to novel context-feature conjunctions. By helping us design hard benchmark tasks that are not solvable by kernel models, our framework grounds research into learning mechanisms that can implement other compositional computations.
What do our findings imply about compositional generalization in more general representations? In this section, we analyzed compositionally structured inputs, as we consider them the best-case scenario for compositional generalization. Accordingly, we expect that the limitations revealed by our theory are also relevant for kernel models with non-compositionally structured representations. We present two pieces of preliminary evidence for this conjecture in the appendix: First, we theoretically prove that randomly sampled representations (that are only compositionally structured in expectation) are also conjunction-wise additive in expectation (Section A.5). Second, we empirically investigate compositional generalization in disentangled representation learning models trained on the DSprites dataset (Locatello et al., 2019), finding that, when averaged across randomly sampled task instances, these models are also limited to conjunction-wise additive computations (Appendix C). These findings suggest that the generalization class we have highlighted in this section captures fundamental limitations of kernel models beyond our theoretical setting.
@@
Conjunction-wise additivity only determines whether kernel models can, in principle, solve a certain task. However, model generalization also depends on whether the model identifies the correct conjunction-wise function. This depends on the model’s representational geometry and the training data statistics. As noted in Section 2, the behavior of kernel models depends on their representation $\phi(x)$ only through the induced kernel $K_{\phi}(x,x^{\prime})$. In this section, we first investigate how different choices in network architecture influence its induced kernel (Section 5.1). Then, we characterize how this kernel impacts task performance on symbolic addition and context dependence (Section 5.2).
@@ 5.1 Overlap salience characterizes compositional representational geometry
To characterize compositionally structured representations, we introduce a new metric: representational salience. We formally define this metric in Section B.1. Intuitively, for $k=1,\dotsc,C$, the salience $S(k;C)$ measures the unique contribution of the subpopulation representing a conjunction of $c$ components. Further, $S(k;C)$ is normalized so that all saliences sum up to one. For example, the multi-hot input only encodes single components and therefore $S(1;C)=1/C$ and, for all $k>1$, $S(k;C)=0$. Note that the model can only use conjunctions whose salience is nonzero. For example, the multi-hot input is constrained to a component-wise sum (as all other saliences are zero).
$S(k;C)$ is computed from the representational similarities $K_{\phi}(x,x^{\prime})=\phi(x)^{T}\phi(x^{\prime})$. This metric is useful because unlike $K_{\phi}(x,x^{\prime})$, it makes immediately apparent how strongly different conjunctions are encoded (Section B.2). Furthermore, it reduces the $C+1$ similarities characterizing the space of compositionally structured representations to $C-1$ free parameters, removing the overall magnitude of the representation (by normalizing) and the baseline activity (by not considering the similarity between inputs with no overlap). This is because these parameters often have negligible impact on model behavior (Section B.3).
@@
We now analyze how $S(k;C)$ evolves over different layers of a random network, using three components as an example (Fig. 3; see Section B.4 for other values of $C$). The input only represents individual components and so $S(2;3)$ and $S(3;3)$ are zero. As the network gets deeper, these saliences increase and $S(1;3)$ decreases. In particular, because $S(3;3)>0$ for all networks with at least one layer, the resulting representations can learn arbitrary training data (using the full conjunction). As depth increases, $S(2;3)$ eventually decreases again, whereas $S(3;3)$ continues to increase.
In fact, we prove that for ReLU networks in the limit of infinite depth, $S(C;C)\to 1$ (Proposition B.2). Put differently, very deep networks exclusively encode the full conjunction. As a result, these models effectively implement a look-up table, memorizing the training data without generalizing to novel combinations of components. In the next section, we find that even a moderately conjunctive representation presents significant challenges to compositional generalization.
@@ 5.2 Kernel models suffer from memorization leak and shortcut bias
While there are many readout weights that can lead to minimal error, kernel models trained with gradient descent learn the weights with minimal $\ell_{2}$-norm (see Section 2). We now characterize how this inductive bias influences compositional generalization. To characterize the full range of compositionally structured representations, we compute the behavior of the model directly using the kernel (see Appendix D). Importantly, our analysis clarifies the behavior of all random weight neural networks, including those covered in the previous section.
Symbolic addition suffers from a memorization leak. We first consider an example task ($z_{1},z_{2}\in\{[-4],\dotsc,[4]\}$; the training set consists of all trials containing a $[0]$, Fig. 2a) and an example representation ($S(1;2)=0.4$). We find that though the model perfectly learns the training cases, it underestimates the test cases by a proportional factor (Fig. 4a). To understand why, we consider the model’s functional form on the training cases: $f(x)=f_{1}(z_{1})+f_{2}(z_{2})+f_{12}(z)$ (see Section 4). Intuitively, $\ell_{2}$-norm minimization tends to yield distributed weights. As a result, unless the salience of the conjunction $S(2;2)$ is zero, the kernel model will associate some non-zero weight with it (i.e. $f_{12}(z_{1},z_{2})\neq 0$). This distorts the test set generalization, $f(x)=f_{1}(z_{1})+f_{2}(z_{2})$.
We call the tendency to use the full conjunction (and its detrimental effect on generalization) “memorization leak.” To characterize it more systematically (and add mathematical rigor to the intuition above), we analytically characterize model behavior on a general family of symbolic addition tasks:
@@
We prove the proposition in Section E.1. It implies that for this entire task family, model generalization is distorted by a proportional factor $m$ (Fig. 4b). The formula implies that $m<1$ as long as $S(1;2)<\tfrac{1}{2}$ (i.e. as long as $S(2;2)>0$). Further, $m$ is smaller for lower $S(1;2)$ and smaller training set size $p$. (Note that $\mathcal{W}$ corresponds to the rows/columns making up the example training sets in Fig. 2a.) Interestingly, these are the only two factors that influence $m$. In particular, even though interpolation is often seen as easier than extrapolation, this does not impact model behavior here.
More broadly, the memorization leak is a ubiquitous issue for statistical learning models trained on compositional tasks: $\ell_{2}$-norm minimization tends to yield distributed weights and so if the conjunctive population is represented, the model will generally end up partially relying on this conjunction. As shown above, this necessarily distorts generalization. This issue also arises in tasks that involve directly decoding specific components, a popular task for evaluating disentangled representations (Locatello et al., 2019; Schott et al., 2022) (see Section E.3 for a minimal example).
@@
Context dependence suffers from a shortcut bias. Next, we empirically analyzed model generalization on context dependence across different representational geometries and training sets. We found that on a given task, each model either generalizes with 100% or 0% accuracy. For CD-3 (the task version leaving out the largest subset of feature combinations), the model only generalizes when $S(2;3)$ is high relative to $S(1;3)$ (Fig. 4c, dark green area). As a result, compositional generalization is highly sensitive to the nonlinearity and depth of the network. In contrast, for CD-2 and CD-1, a much wider range of representational geometries generalizes successfully.
To understand this phenomenon, we determined the total magnitude of model weights associated with the different conjunctions (Fig. 4d). We found that unsuccessful models (e.g. the blue color in Fig. 4d) had much larger magnitudes associated with the context component and the full conjunction. Notably, on CD-3, context is highly correlated with the target and can predict $\tfrac{2}{3}$ of the training data. Models with high $S(1;3)$ (context) and $S(3;3)$ (full conjunction) exploit this context shortcut and use the full conjunction to learn the remaining training data. This strategy explains why these models fail to generalize to the test set. For CD-2, context is much less predictive on the training data (accuracy of $\tfrac{9}{16}$). This explains why only very low $S(2;3)$ yields failure to generalize on CD-2 or CD-1.
Our analysis shows how model and task structure interact to either give rise to a generalizable rule or a statistical shortcut. This highlights that for compositional tasks with strong spurious correlations, model behavior will be highly sensitive to architectural hyperparameters affecting the representational geometry such as depth and nonlinearity. Indeed, this sensitivity to minor experimental choices could explain why the literature has been so divided on the usefulness of disentangled representations.
6 Our theory can describe the behavior of deep networks on conjunction-wise additive tasks